# Intensified optical camera with Timepix4 readout

### v1

@@ 4 Conclusions
Table 1 presents a comparison of timing resolutions (RMS) achieved in this study for different configurations of Timepix4 camera. The resolution deteriorates after addition of the intensifier mostly due to appearance of a scintillator as an intermediate step in the amplification chain. The P47 scintillator typically used in fast image intensifiers has a rise time of about 7 ns and, therefore, can limit the time resolution in situations with insufficient gain. Faster scintillators are becoming available [51, 55], which can be used in intensifiers and also for similar MCP-based amplification schemes employed for ion and electron detection.
@@
| intensifier: mild ToT selection + timewalk correction | 1.42 | large detection efficiency |
| intensifier: high ToT selection | 0.55 | small detection efficiency |
In summary, we presented the first characterization of an optical camera based on the Timepix4 chip coupled to an optical silicon sensor and image intensifier, demonstrating its potential for sub-nanosecond scale, time-resolved single-photon imaging. The system achieved time resolutions down to about 0.3 ns without the intensifier and as low as 0.55 ns with intensifier in the single-photon regime and selection of signals with high amplitude. For a more inclusive case of amplitude selections covering 82% of events the resolution with intensifier is equal to 1.4 ns, which is still a factor of two to three better than the resolution achieved in the past with the Timepix3 cameras [3, 2]. These measurements validate the capability of the Timepix4-based optical cameras to perform fast, high-resolution optical time-stamping. The observed performance marks a significant improvement over previous Timepix3-based systems, particularly in terms of timing precision, pixel count and readout rate.
@@ Acknowledgements.
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and Czech Ministry of Education, Youth and Sports Project No. LM2023040 CERN-CZ. We thank Lou-Ann Pestana De Sousa for help with preparation of manuscript.References
- [1] Fisher-Levine, M. and Nomerotski, A., “TimepixCam: a fast optical imager with time-stamping,” Journal of Instrumentation 11(03), C03016 (2016).
- [2] Nomerotski, A., Chekhlov, M., Dolzhenko, D., Glazenborg, R., Farella, B., Keach, M., Mahon, R., Orlov, D., and Svihra, P., “Intensified Tpx3Cam, a fast data-driven optical camera with nanosecond timing resolution for single photon detection in quantum applications,” Journal of Instrumentation 18, C01023 (Jan. 2023).
- [3] Zhao, A., van Beuzekom, M., Bouwens, B., Byelov, D., Chakaberia, I., Cheng, C., Maddox, E., Nomerotski, A., Svihra, P., Visser, J., Vrba, V., and Weinacht, T., “Coincidence velocity map imaging using tpx3cam, a time stamping optical camera with 1.5 ns timing resolution,” Review of Scientific Instruments 88 (Nov. 2017).
- [4] Albrechtsen, S. H., Schouder, C. A., Viñas Muñoz, A., Christensen, J. K., Engelbrecht Petersen, C., Pi, M., Barranco, M., and Stapelfeldt, H., “Observing the primary steps of ion solvation in helium droplets,” Nature 623, 319–323 (Nov. 2023).
- [5] Sandström, E., Huysmans, P., Giskes, F., Laeven, P., Van Nuffel, S., Heeren, R. M. A., and Anthony, I. G. M., “Improvements in fast mass microscopy for large-area samples,” Analytical Chemistry 96, 18037–18042 (Oct. 2024).
- [6] Bromberger, H., Passow, C., Pennicard, D., Boll, R., Correa, J., He, L., Johny, M., Papadopoulou, C. C., Tul-Noor, A., Wiese, J., Trippel, S., Erk, B., and Küpper, J., “Shot-by-shot 250 khz 3d ion and mhz photoelectron imaging using timepix3,” Journal of Physics B: Atomic, Molecular and Optical Physics 55, 144001 (June 2022).
- [7] Roberts, A., Svihra, P., Al-Refaie, A., Graafsma, H., Küpper, J., Majumdar, K., Mavrokoridis, K., Nomerotski, A., Pennicard, D., Philippou, B., Trippel, S., Touramanis, C., and Vann, J., “First demonstration of 3d optical readout of a TPC using a single photon sensitive timepix3 based camera,” Journal of Instrumentation 14, P06001–P06001 (June 2019).
- [8] Hirvonen, L. M., Fisher-Levine, M., Suhling, K., and Nomerotski, A., “Photon counting phosphorescence lifetime imaging with timepixcam,” Review of Scientific Instruments 88(1), 013104 (2017).
- [9] Sen, R., Hirvonen, L. M., Zhdanov, A., Svihra, P., Andersson-Engels, S., Nomerotski, A., and Papkovsky, D., “New luminescence lifetime macro-imager based on a tpx3cam optical camera,” Biomed. Opt. Express 11, 77–88 (Jan 2020).
- [10] Sen, R., Zhdanov, A. V., Bastiaanssen, T. F. S., Hirvonen, L. M., Svihra, P., Fitzgerald, P., Cryan, J. F., Andersson-Engels, S., Nomerotski, A., and Papkovsky, D. B., “Mapping o2 concentration in ex-vivo tissue samples on a fast PLIM macro-imager,” Scientific Reports 10 (Nov. 2020).
- [11] D'Amen, G., Keach, M., Nomerotski, A., Svihra, P., and Tricoli, A., “Novel imaging technique for -particles using a fast optical camera,” Journal of Instrumentation 16, P02006–P02006 (Feb. 2021).
- [12] Losko, A. S., Han, Y., Schillinger, B., Tartaglione, A., Morgano, M., Strobl, M., Long, J., Tremsin, A. S., and Schulz, M., “New perspectives for neutron imaging through advanced event-mode data acquisition,” Scientific Reports 11(1) (2021).
- [13] Yang, J., Zhou, J., Jiang, X., Tan, J., Zhang, L., Zhou, J., Zhou, X., Yang, W., Xia, Y., Chen, J., Sun, X., Zhang, Q., Li, J., Sun, Z., and Chen, Y., “A novel energy resolved neutron imaging detector based on a time stamping optical camera for the CSNS,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 1000, 165222 (June 2021).
- [14] Wolfertz, A., Gustschin, A., Schulz, M., Long, A. M., Khaplanov, A., Hirsh, T. Y., Nomerotski, A., Morgano, M., Tremsin, A., Mauri, G., Sykora, G. J., and Losko, A., “Lumacam: a novel class of position-sensitive event mode particle detectors using scintillator screens,” Scientific Reports 14 (Dec. 2024).
- [15] Ianzano, C., Svihra, P., Flament, M., Hardy, A., Cui, G., Nomerotski, A., and Figueroa, E., “Fast camera spatial characterization of photonic polarization entanglement,” Scientific Reports 10 (Apr 2020).
- [16] Zhang, Y., England, D., Nomerotski, A., Svihra, P., Ferrante, S., Hockett, P., and Sussman, B., “Multidimensional quantum-enhanced target detection via spectrotemporal-correlation measurements,” Phys. Rev. A 101, 053808 (May 2020).
- [17] Svihra, P., Zhang, Y., Hockett, P., Ferrante, S., Sussman, B., England, D., and Nomerotski, A., “Multivariate discrimination in quantum target detection,” arXiv:2005.00612 (May 2020).
- [18] Nomerotski, A., Keach, M., Stankus, P., Svihra, P., and Vintskevich, S., “Counting of Hong-Ou-Mandel Bunched Optical Photons Using a Fast Pixel Camera,” Sensors 20(12) (2020).
- [19] Zhang, Y., England, D., Nomerotski, A., and Sussman, B., “High speed imaging of spectral-temporal correlations in hong-ou-mandel interference,” Optics Express 29, 28217 (Aug. 2021).
- [20] Gao, X., Zhang, Y., D’Errico, A., Heshami, K., and Karimi, E., “High-speed imaging of spatiotemporal correlations in hong-ou-mandel interference,” Optics Express 30, 19456 (May 2022).
- [21] Zhang, Y., Orth, A., England, D., and Sussman, B., “Ray tracing with quantum correlated photons to image a three-dimensional scene,” Physical Review A 105 (Jan. 2022).
- [22] Zhukas, L. A., Svihra, P., Nomerotski, A., and Blinov, B. B., “High-fidelity simultaneous detection of a trapped-ion qubit register,” Physical Review A 103 (June 2021).
- [23] Zhukas, L. A., Millican, M. J., Svihra, P., Nomerotski, A., and Blinov, B. B., “Direct observation of ion micromotion in a linear paul trap,” Physical Review A 103 (Feb. 2021).
- [24] Kato, A., Goel, A., Lee, R., Ye, Z., Karki, S., Liu, J. J., Nomerotski, A., and Blinov, B. B., “Two-tone doppler cooling of radial two-dimensional crystals in a radio-frequency ion trap,” Physical Review A 105 (Feb. 2022).
- [25] Vallerga, J., Tremsin, A., DeFazio, J., Michel, T., Alozy, J., Tick, T., and Campbell, M., “Optical mcp image tube with a quad timepix readout: initial performance characterization,” Journal of Instrumentation 9(05), C05055 (2014).
- [26] Bolzonella, R., Alozy, J., Ballabriga, R., Biesuz, N., Campbell, M., Cavallini, V., Ramusino, A. C., Fiorini, M., Franzoso, E., Guarise, M., Cudie, X. L., Romolini, G., and Saputi, A., “Development and characterization of hybrid mcp-pmt with embedded timepix4 asic used as pixelated anode,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 1082, 170965 (Feb. 2026).
- [27] Jachura, M. and Chrapkiewicz, R., “Shot-by-shot imaging of Hong–Ou–Mandel interference with an intensified sCMOS camera,” Optics letters 40(7), 1540–1543 (2015).
- [28] Jost, B. M., Sergienko, A. V., Abouraddy, A. F., Saleh, B. E., and Teich, M. C., “Spatial correlations of spontaneously down-converted photon pairs detected with a single-photon-sensitive CCD camera,” Optics Express 3(2), 81–88 (1998).
- [29] Brida, G., Caspani, L., Gatti, A., Genovese, M., Meda, A., and Berchera, I. R., “Measurement of Sub-Shot-Noise Spatial Correlations without Background Subtraction,” Phys. Rev. Lett. 102, 213602 (May 2009).
- [30] Zhang, L., Neves, L., Lundeen, J. S., and Walmsley, I. A., “A characterization of the single-photon sensitivity of an electron multiplying charge-coupled device,” Journal of Physics B: Atomic, Molecular and Optical Physics 42(11), 114011 (2009).
- [31] Fickler, R., Krenn, M., Lapkiewicz, R., Ramelow, S., and Zeilinger, A., “Real-time imaging of quantum entanglement,” Scientific reports 3(1), 1914 (2013).
- [32] Avella, A., Ruo-Berchera, I., Degiovanni, I. P., Brida, G., and Genovese, M., “Absolute calibration of an EMCCD camera by quantum correlation, linking photon counting to the analog regime,” Optics letters 41(8), 1841–1844 (2016).
- [33] Reichert, M., Sun, X., and Fleischer, J. W., “Quality of spatial entanglement propagation,” Physical Review A 95(6), 063836 (2017).
- [34] Moreau, P.-A., Toninelli, E., Gregory, T., and Padgett, M. J., “Imaging with quantum states of light,” Nature Reviews Physics 1(6), 367–380 (2019).
- [35] Vallerga, J., Tremsin, A., DeFazio, J., Michel, T., Alozy, J., Tick, T., and Campbell, M., “Optical mcp image tube with a quad timepix readout: initial performance characterization,” Journal of Instrumentation 9, C05055–C05055 (May 2014).
- [36] Alozy, J., Biesuz, N., Campbell, M., Cavallini, V., Cotta Ramusino, A., Fiorini, M., Guarise, M., and Llopart Cudie, X., “Development of a single-photon imaging detector with pixelated anode and integrated digital read-out,” Journal of Instrumentation 17, C06007 (June 2022).
- [37] Karl, S., Leopold, V., Richter, S., Prokazov, Y., Turbin, E., Sintotskiy, G., Orlov, D., and von Zanthier, J., “High-throughput single photon detection for effective stellar intensity interferometry,” in Advanced Photon Counting Techniques XIX , Itzler, M. A., McIntosh, K. A., and Bienfang, J. C., eds., 2, SPIE (May 2025).
- [38] Charbon, E., “Single-photon imaging in complementary metal oxide semiconductor processes,” Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering Sciences 372(2012), 20130100 (2014).
- [39] Perenzoni, M., Pancheri, L., and Stoppa, D., “Compact SPAD-based pixel architectures for time-resolved image sensors,” Sensors 16(5), 745 (2016).
- [40] Gasparini, L., Bessire, B., Unternährer, M., Stefanov, A., Boiko, D., Perenzoni, M., and Stoppa, D., “SUPERTWIN: towards 100kpixel CMOS quantum image sensors for quantum optics applications,” in Quantum Sensing and Nano Electronics and Photonics XIV , 10111, 404–414, SPIE (2017).
- [41] Bruschini, C., Homulle, H., Antolovic, I. M., Burri, S., and Charbon, E., “Single-photon avalanche diode imagers in biophotonics: review and outlook,” Light: Science & Applications 8(1), 87 (2019).
- [42] Morimoto, K., Ardelean, A., Wu, M.-L., Ulku, A. C., Antolovic, I. M., Bruschini, C., and Charbon, E., “Megapixel time-gated SPAD image sensor for 2D and 3D imaging applications,” Optica 7(4), 346–354 (2020).
- [43] Lubin, G., Tenne, R., Ulku, A. C., Antolovic, I. M., Burri, S., Karg, S., Yallapragada, V. J., Bruschini, C., Charbon, E., and Oron, D., “Heralded spectroscopy reveals exciton–exciton correlations in single colloidal quantum dots,” Nano letters 21(16), 6756–6763 (2021).
- [44] Wojtkiewicz, M., Rae, B., and Henderson, R. K., “Review of Back-Side Illuminated 3-D-Stacked SPADs for Time-of-Flight and Single-Photon Imaging,” IEEE Transactions on Electron Devices (2024).
- [45] Hadfield, R. H., “Single-photon detectors for optical quantum information applications,” Nature Photonics 3, 696–705 (Dec. 2009).
- [46] Seitz, P. and Theuwissen, A. J., Single-photon imaging , vol. 160, Springer Science & Business Media (2011).
- [47] Llopart, X., Alozy, J., Ballabriga, R., Campbell, M., Casanova, R., Gromov, V., Heijne, E., Poikela, T., Santin, E., Sriskaran, V., Tlustos, L., and Vitkovskiy, A., “Timepix4, a large area pixel detector readout chip which can be tiled on 4 sides providing sub-200 ps timestamp binning,” Journal of Instrumentation 17, C01044 (Jan. 2022).
- [48] Heijhoff, K., Akiba, K., Ballabriga, R., van Beuzekom, M., Campbell, M., Colijn, A., Fransen, M., Geertsema, R., Gromov, V., and Llopart Cudie, X., “Timing performance of the timepix4 front-end,” Journal of Instrumentation 17, P07006 (July 2022).
- [49] Poikela, T., Plosila, J., Westerlund, T., Campbell, M., Gaspari, M. D., Llopart, X., Gromov, V., Kluit, R., Beuzekom, M. v., Zappon, F., Zivkovic, V., Brezina, C., Desch, K., Fu, Y., and Kruth, A., “Timepix3: a 65k channel hybrid pixel readout chip with simultaneous toa/tot and sparse readout,” Journal of Instrumentation 9, C05013–C05013 (May 2014).
- [50] Nomerotski, A., Chakaberia, I., Fisher-Levine, M., Janoska, Z., Takacs, P., and Tsang, T., “Characterization of timepixcam, a fast imager for the time-stamping of optical photons,” Journal of instrumentation 12(01), C01017 (2017).
- [51] Winter, B., King, S. J., Brouard, M., and Vallance, C., “A fast microchannel plate-scintillator detector for velocity map imaging and imaging mass spectrometry,” Review of Scientific Instruments 85 (Feb. 2014).
- [52] Tsigaridas, S., Beuzekom, M., Graaf, H., Hartjes, F., Heijhoff, K., Hessey, N., de Jong, P., and Prodanovic, V., “Timewalk correction for the timepix3 chip obtained with real particle data,” Nuclear Instruments and Methods in Physics Research Section A: Accelerators, Spectrometers, Detectors and Associated Equipment 930, 185–190 (June 2019).
- [53] Bolzonella, R., Alozy, J., Ballabriga, R., van Beuzekom, M., Biesuz, N., Campbell, M., Cardarelli, P., Cavallini, V., Coco, V., Cotta Ramusino, A., Fiorini, M., Gromov, V., Guarise, M., Llopart Cudie, X., Okamura, S., Romolini, G., Saputi, A., and Vitkovskiy, A., “Timing resolution performance of timepix4 bump-bonded assemblies,” Journal of Instrumentation 19, P07021 (July 2024).
- [54] Dorkel, J. and Leturcq, P., “Carrier mobilities in silicon semi-empirically related to temperature, doping and injection level,” Solid-State Electronics 24, 821–825 (Sept. 1981).
- [55] Zapadlík, O., Nikl, M., Polák, J., Průša, P., and Linhart, V., “Engineering of yag:ce to improve its scintillation properties,” Optical Materials: X 15, 100165 (Aug. 2022).

### v2

@@ 4 Conclusions
Table 1 presents a comparison of timing resolutions (RMS) achieved in this study for different configurations of the Timepix4 camera. The resolution deteriorates after the addition of the intensifier, mostly due to the appearance of a scintillator as an intermediate step in the amplification chain. The P47 scintillator typically used in fast image intensifiers has a rise time of about 7 ns and, therefore, can limit the time resolution in situations with insufficient gain. Faster scintillators are becoming available [53, 57], which can be used in intensifiers and also for similar MCP-based amplification schemes employed for ion and electron detection. Another opportunity to increase the signal amplitude with the aim of improving time resolution is to employ optical sensors based on low-gain avalanche diodes (LGAD), which typically provide a gain of 5 to 10 [58, 59, 60]. Work is in progress to implement and evaluate these options.
@@
| intensifier: mild ToT selection + timewalk correction | 1.0–1.5 | large detection efficiency |
| intensifier: high ToT selection | 0.6–0.7 | small detection efficiency |
In summary, we presented the first characterization of an optical camera based on the Timepix4 chip coupled to an optical silicon sensor and image intensifier, demonstrating its potential for sub-nanosecond scale, time-resolved single-photon imaging. The system achieved time resolutions down to about 0.3 ns without the intensifier and as low as 0.6 ns with the intensifier in the single-photon regime and selection of signals with high amplitude. For two more inclusive cases of amplitude selections covering, respectively, 43% and 83% of events, the resolutions with the intensifier are equal to 1.0 ns and to 1.5 ns, which is still a factor of two to three better than the resolution achieved in the past with Timepix3 cameras [3, 2]. These measurements validate the capability of the Timepix4-based optical cameras to perform fast, high-resolution optical time-stamping. The observed performance marks a significant improvement over previous Timepix3-based systems, particularly in terms of timing precision, pixel count, and readout rate.
@@ Acknowledgments
This research was supported by the Czech Science Foundation (GACR) under Project No. 25-15534M and Czech Ministry of Education, Youth and Sports Project No. LM2023040 CERN-CZ. We thank Lou-Ann Pestana De Sousa for help with preparation of manuscript.